OpenStudy (anonymous):

Solve the system of equations using matrices. Use Gaussian elimination with back-substitution.

OpenStudy (anonymous):

x + y + z = -5 x - y + 3z = -1 4x + y + z = -2 Can someone show me how to solve this? step by step please

OpenStudy (abb0t):

Ok. I'm going to assume this is for Linear Algebra and you're familiar with the notation for the operations, correct? So I am not going to draw it out because it will take me too long.

OpenStudy (anonymous):

This is pre-cal. & Im not sure?